Circular No. 8364

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004cy IN ESO 403-G9 Y. Chen, Z. Lin, K. Huang, H. Lin, and W. Ip, Institute of Astron-

omy, National Central University, Tawain; and Y. Qiu and J. Hu, National Astronomical Observatories of China, report the discovery of an apparent supernova (mag ∼ 15.6) on unfiltered CCD images taken on Apr. 5.87 UT with the 1-m Lulin Observatory telescope. SN 2004cy is located at α = 21^{h}34^{m}25^{s}.67, δ = −33^{o}23′13′′.7 (equinox 2000.0), which is 25′′.7 west and 6′′.1 south of the center of ESO 403-G9. Approximate unfiltered magnitudes for the new object: 2003 Oct. 31, [19.0; 2004 Apr. 25.83, 17.5; May 27.77, 17.8; 28.77, 17.9; June 9.80, 18.2; 25.77, 18.2; 26.77, 18.4; 27.74, 18.3.

SUPERNOVA 2004cv IN MCG +03-41-120 G. Schwarz and A. Karam, Steward Observatory (SO); R. M. Wagner,

LBT Observatory; and J. G. Rohrbach and S. Starrfield, Arizona State University, report that a spectrum (range 420–740 nm; resolution 0.65 nm) of SN 2004cv (cf. IAUC 8362), obtained on June 26.22 UT with the SO 2.3m Bok telescope (+ Boller & Chivens spectrograph), shows it to be type-Ia supernova near maximum light. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 11018 km/s, the supernova expansion velocity is 13700 km/s for the Si II feature (rest 635.5 nm).

C. L. Gerardy, University of Texas at Austin, reports that a lowresolution optical spectrogram (resolution 300; range 500–1000 nm) of SN 2004cv, obtained on June 29.28 UT with the 9.2-m Hobby/Eberly Telescope (+ Marcario Low-Resolution Spectrograph) by B. Roman and F. Deglman, indicates it to be a type-Ia supernova near or slightly before maximum light. Adopting the redshift of MCG +03-41-120 (11018 km/s) from the NED, the expansion velocity of the Si II feature (rest wavelength 635.5 nm) is 12500 km/s.

SUPERNOVA 2004cx IN NGC 7755 M. Salvo and B. Schmidt, Australian National University (ANU); and

G. Altavilla, University of Barcelona, on behalf of the ERTN (cf. IAUC 7961), report that inspection of a poor-S/N spectrum (range 590–680 nm; resolution 6000) of SN 2004cx (cf. IAUC 8362), taken by L. Kiss and S. Meszaros (University of Sydney) with the ANU 2.3-m telescope (+ Double Beam Spectrograph) at Siding Spring on June 30.80 UT, shows the profile of a young type-II supernova. The expansion velocity, measured from the minimum of the P-Cyg profile of the Hα line, is around 5000 km/s.

2004 July 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green